OpenStudy (anonymous):

9x4-16y2 formula: a3-b3= (a-b) (a2+ab+b2)

OpenStudy (anonymous):

The f-la is a² – b²= ( a + b ) ( a – b ) (3x^2)^2-(4y)^2=(3x^2-4y)(3x^2+4y)

OpenStudy (anonymous):

im factoring cubics..its different

OpenStudy (anonymous):

But you don't have cubes there